### v1

# Is In-Context Learning Sufficient for Instruction Following in LLMs?

Abstract
In-context learning (ICL) allows LLMs to learn from examples without changing their weights, which is a particularly promising capability for long-context LLMs that can potentially learn from many examples. Recently, Lin et al. (2024) proposed Urial, a method using only three in-context examples to align base LLMs, achieving non-trivial instruction following performance. In this work, we show that, while effective, ICL alignment with Urial still underperforms compared to instruction fine-tuning on established benchmarks such as MT-Bench and AlpacaEval 2.0 (LC), especially with more capable base LMs. Unlike for tasks such as classification, translation, or summarization, adding more ICL demonstrations for long-context LLMs does not systematically improve instruction following performance. To address this limitation, we derive a greedy selection approach for ICL examples that noticeably improves performance, yet without bridging the gap to instruction fine-tuning. Finally, we provide a series of ablation studies to better understand the reasons behind the remaining gap, and we show how some aspects of ICL depart from the existing knowledge and are specific to the instruction tuning setting. Overall, our work advances the understanding of ICL as an alignment technique. We provide our code at https://github.com/tml-epfl/icl-alignment.
Machine Learning, ICML
1 Introduction
The large-scale pre-training phase allows Large Language Models (LLMs) to acquire extensive knowledge and capabilities. However, these base models struggle to follow instructions directly from prompts, necessitating further fine-tuning to be used as conversational models. Traditional alignment methods include Supervised Fine-Tuning (SFT) on instruction datasets (IFT, Chung et al., 2022) or preference data (DPO and KTO, Rafailov et al., 2023; Ethayarajh et al., 2024), and reinforcement learning (RLHF, Ouyang et al., 2022). Inspired by Brown et al. (2020) who showed that LLMs can learn from demonstrations provided as part of the input—a concept known as in-context learning (ICL)—Lin et al. (2024) have recently studied the feasibility of in-context alignment. They found that including merely three carefully selected question-answer pairs in the prompt is sufficient to make base models follow instructions and interact with users at a similar level to instruction-finetuned models on their own benchmark. The possibility of customizing base models via ICL without changing weights is potentially game-changing due to its simplicity and flexibility.
In this work, we evaluate the Urial prompt strategy proposed by Lin et al. (2024) across several base models, including GPT-4-Base [^1] We received access to the base GPT-4 model via the OpenAI Researcher Access Program., on established instruction following benchmarks like MT-Bench (Zheng et al., 2023) and AlpacaEval 2.0 (Li et al., 2023). Although Urial achieves reasonable performance, it still lags behind instruction-finetuned models. In response, we test various strategies to improve Urial, particularly leveraging models with extensive context windows which allow for longer in-context prompts. First, we study the effect of many-shot in-context learning by adding high-quality demonstrations from existing instruction datasets. This approach, however, leads only to limited improvements compared to Urial, without a clear improvement trend for a large number of in-context examples. This is in contrast to many-shot ICL for tasks like summarization, translation, or classification, where providing many examples is clearly beneficial (Agarwal et al., 2024; Bertsch et al., 2024). Then we propose a simple greedy algorithm to select the in-context examples which optimize the MT-Bench score. This selection scheme allows us to nearly close the gap between Mistral-7b-v0.2 and Mistral-7b-Instruct-v0.2 by adding two additional demonstrations to the Urial prompt. Finally, we revisit the findings of Min et al. (2022) regarding what makes ICL effective: unlike for classification and multiple choice tasks, providing questions with correct answers is crucial for the effectiveness of ICL on instruction-following tasks. Furthermore, adding incorrect demonstrations further degrades the model performance. Overall, our work provides a more nuanced picture of ICL as an alignment technique compared to previous works.
2 Systematic evaluation of Urial
First, we compare the performance of several base models with the Urial in-context prompt to that of their instruction fine-tuned versions. Table 1 shows the results on MT-Bench (Zheng et al., 2023) which is one of the most popular benchmarks for instruction following ability of LLMs. We note, however, that Lin et al. (2024) originally tested Urial on their own dataset, in which only $8\%$ of the examples are obtained from MT-Bench. We compare several LLMs of different sizes and capabilities, ranging from Llama-2-7B (Touvron et al., 2023) to the base GPT-4 model (OpenAI, 2023). For simplicity, we consider the official instruction-tuned LLMs, i.e., provided from the same source of the base models, even if there exist several third-party fine-tuned versions which can achieve higher scores. We observe that base models with Urial achieve competitive performance on the 1st-turn score, but cannot match the performance of their instruction fine-tuned counterparts in all cases except for Llama-2-70B and Mistral-7B-v0.1 (note that both were originally included in Lin et al. (2024), unlike most others). Conversely, the 2nd-turn score with Urial is significantly worse than for instruction-tuned LLMs: we hypothesize that this is because no multi-turn demonstrations are given in context. Indeed, Zhou et al. (2023) show that, even in IFT, having a few multi-turn training examples significantly improve the performance of the model on multi-turn conversations. Because of this, in the rest of the paper, we focus on the 1st-turn score and single-round conversations, and leave extending ICL for multi-round conversations to future work. We detail the breakdown of the results from Table 1 over categories in Table 5.
Table 1: Systematic comparison of Urial to finetuned models on MT-Bench across different base LLMs. * denotes the result taken from the Urial GitHub repository.
Model 1st-turn 2nd-turn Average
Llama-2-7B + Urial * $5.75$ $3.91$ $4.83$
Llama-2-7B-Instruct 7.14 5.91 6.53
Llama-2-70B + Urial * 7.61 $6.61$ $7.11$
Llama-2-70B-Instruct $7.37$ 7.03 7.20
Llama-3-8B + Urial * $6.84$ $4.65$ $5.75$
Llama-3-8B-Instruct 8.29 7.42 7.86
Llama-3-70B + Urial * $7.71$ $5.09$ $6.40$
Llama-3-70B-Instruct 8.96 8.51 8.74
Mistral-7B-v0.1 + Urial * 7.49 $5.86$ $6.67$
Mistral-7B-Instruct-v0.1 $7.31$ 6.39 6.85
Mistral-7B-v0.2 + Urial * $6.99$ $5.55$ $6.27$
Mistral-7B-Instruct-v0.2 8.06 7.21 7.64
Mixtral-8x22B-v0.1-4bit + Urial $8.28$ $7.14$ $7.71$
Mixtral-8x22B-Instruct-v0.1-4bit 8.78 8.25 8.52
GPT-4-Base + Urial $7.96$ $5.04$ $6.50$
GPT-4 (March 2023) * 8.96 9.03 8.99
3 Many-shot ICL for instruction following
In the following, we explore whether we can close the gap between ICL and fine-tuning on instruction following tasks. In particular, we focus on Mistral-7B-v0.2 (Jiang et al., 2023) since (a) both Urial and the aligned model achieve competitive performance in Table 1, and (b) this base model has a 32k context window which allows us to use many ICL examples. Moreover, we measure performance with both the MT-Bench 1st-turn score and the AlpacaEval 2.0 win rate, and report the results in Table 2. More details can be found in App. A.
Figure 1: The results of scaling experiments with different sampling strategies on Mistral-7B-v0.2-32k. The black dashed line refers to the 1st-turn MT-Bench score of Mistral-7B-Instruct-v0.2. ICL with more demonstrations does not bridge the gap between the base model and its aligned counterpart. Surprisingly, with only three examples from Urial and questions without answers, the base model always achieves better instruction-following performance than with the full question-answer pairings.
Experimenting with different selection criteria. First, we test adding random demonstrations (after simple pre-filtering) to the three already included in Urial. For this, we score the quality of the examples in Evol-Instruct-70k (Xu et al., 2024) with GPT-3.5-Turbo, and randomly sample demonstrations among the highest quality instructions (those with score at least 4.5 on a 0-5 scale, see Fig. 4). Table 2 shows that adding 17 or 37 random examples (which makes 20 and 40 examples in total) provides only moderate improvements, around $+0.3$, and Fig. 1 indicates that such an increase quickly reaches a plateau (see the Evol-Instruct-Random curve), without clear benefits from scaling up to 50 demonstrations. We also test a more refined strategy: we sample instructions from OpenHermes (Teknium, 2023), which includes annotations on the topic of each example, to ensure diversity between categories. This approach is however worse than Urial, regardless of the number of ICL examples used (see the OpenHermes-Diversity curve in Fig. 1). Finally, inspired by the unsupervised ICL of Agarwal et al. (2024), we re-use the OpenHermes examples, this time without the answers (questions only) and adding instead the three Urial complete demonstrations (both questions and answers). This strategy significantly improves the MT-Bench score (from 7.0 to 7.5, see the OpenHermes-Unsupervised curve) and is slightly better than using the random examples from Evol-Instruct.
Figure 2: The distribution of the 1st-turn MT-Bench score (GPT-4-Turbo as judge) on Mistral-7B-v0.2-32k obtained by adding 100 instructions from Evol-Instruct-70k as a 4th demonstration to Urial. The black dashed line refers to the 1st-turn MT-Bench score of the base model with Urial. A majority of the 4th examples have positive contributions to the model’s instruction-following performance.
Adding examples based on greedy search. Next, we try to greedily maximize the MT-Bench score by testing 100 high-quality instructions from Evol-Instruct-70k as the 4th additional example to Urial. For each resulting ICL prompt, we compute the MT-Bench score with GPT-4-Turbo instead of GPT-4 as the former is faster, cheaper, and mitigates potential overfitting to the exact benchmark score. In most cases, there is an improvement, although it is often small (see the score distribution in Fig. 2). Computing the true MT-Bench (GPT-4 as judge) with the best performing point as the 4th ICL example yields a significant improvement over Urial, from 6.99 to 7.46 (Table 2). We then repeat this procedure sequentially to find a 5th and a 6th demonstration while restricting the search space at each round (by keeping only the instructions leading to a high enough MT-Bench score) to reduce the computational cost. As shown in Table 2, this strategy achieves 7.69 MT-Bench score, almost closing the gap with Mistral-7B-Instruct-v0.2 (8.06). In addition, it even improves AlpacaEval 2.0 win rate from 8.09 of Urial to 9.22 without directly optimizing for it. This improvement indicates that the found IC prompt does not completely overfit to the MT-Bench score. Finally, we use the scores obtained in the first round of the greedy search to scale-up the number of ICL examples: the Evol-Instruct-Greedy curve in Fig. 1 is obtained by progressively adding the instructions with the highest scores. However, adding demonstrations in this way beyond the 6 examples does not consistently improve the MT-Bench score.
Takeaways. These results show that, unlike in the setting of Agarwal et al. (2024); Bertsch et al. (2024), simply scaling the number of ICL examples is not sufficient to consistently improve the instruction following performance. However, a greedy search approach can help to bridge the gap to the results of fine-tuned models. We highlight that our greedy search is performed with a limited computational cost, i.e., the set of candidate examples is small and the judge is slightly suboptimal (GPT-4-Turbo instead of GPT-4). We expect that more effective ICL prompts could be found with additional resources.
Table 2: Improved in-context demonstrations for Mistral-7B-v0.2-32k base model. We report the 1st-turn score on MT-Bench and the length-controlled (LC) win rates on AlpacaEval 2.0 when using different IC prompts.
Model MT-Bench (1st) AlpacaEval 2.0
Urial (3 examples) $6.99$ $8.09$
Urial + 17 random examples $7.22$ $8.43$
Urial + 37 random examples $7.33$ $7.18$
Urial + greedy search (1 ex.) $7.46$ $7.91$
Urial + greedy search (2 ex.) 7.69 $8.38$
Urial + greedy search (3 ex.) $7.68$ $9.22$
4 Understanding ICL for instruction following
In the following, we provide further analyses of the challenges associated with ICL in the context of alignment.
4.1 Importance of question-answer matching for in-context demonstrations
Min et al. (2022) showed that using demonstrations with random labels does not significantly impair the results of ICL on classification and multiple choice tasks. We conduct a similar study for instruction following and report the results of the 1st-turn MT-Bench score in Table 3. Denoting $\{(X_{i},Y_{i})\}_{i}$ the set of in-context examples, with $X_{i}$ the query and $Y_{i}$ the corresponding ground-truth answer, we test several configurations on two sets of $\{(X_{i},Y_{i})\}_{i}$: the 3 Urial examples and the 3 Urial examples with the 3 examples found by our greedy search from Sec. 3. First, we do not use any demonstration, i.e., we directly use the original base model (Mistral-7B-v0.2-32k). As expected, this configuration performs poorly (2.93). Second, providing the questions without the answers leads to surprisingly good results ($\approx$6 MT-Bench score). Third, we permute the answers $Y_{i}$s with a circular shift of one position, so that all correct answers are still contained in the prompt but matched with the wrong question. This procedure significantly degrades the performance, especially with more examples (Urial + Greedy Search) where only a score of 1.80 is obtained (note that the minimum score is 1). Next, for each question, we sample a new answer from those provided for other instructions in the same category (in-domain). For example, a question about coding is paired with an answer from a different coding question. This approach ensures that while the content of the answer may be incorrect, its style and format remain appropriate. Although it is worse than the original one, this configuration achieves quite a good score (6.24 and 5.35). Finally, we sample answers from those of instructions from different, out-of-domain categories, so that even their style is incorrect compared to what implied by each question: this sampling leads to the worst instruction following results.
These results show that not all the conclusions from Min et al. (2022) apply to ICL used for instruction following. Most importantly, using answers with correct content and, especially, correct style is crucial for the success of ICL. This property becomes even more evident when increasing the number of examples (to 6 instead of 3). In that case, not using the correct answers significantly degrades the performance. In turn, this result suggests that ICL for instruction following works differently compared to other less open-ended tasks. This intricacy might explain the gap between ICL and fine-tuning illustrated in Sec. 3.
Table 3: Importance of question-answer matching for in-context demonstrations. We report the 1st-turn MT-Bench score of Mistral-7B-v0.2-32k when varying the structure of the in-context examples $\{(X_{i},Y_{i})\}_{i}$, where $X_{i}$ is the query and $Y_{i}$ is the corresponding ground-truth answer.
In-context prompt Urial               (3 examples) Urial + Greedy Search (6 examples)
no demonstrations 2.93 2.93
$X$ only (no $Y$) 5.90 5.97
circular shift of $Y$ 5.04 1.80
in-domain random $Y$ 6.24 5.35
out-of-domain random $Y$ 4.13 1.74
correct $Y$ 6.99 7.68
4.2 Transferability of in-context examples
In Table 4, we report the performance of applying the in-context examples found by greedy search (added to Urial) on Mistral-7B-v0.2-32k (see Sec. 3) to Llama-2-7B-80k (Fu et al., 2024) and Mixtral-8x22B-v0.1-4bit (Jiang et al., 2024). Adding the new examples does not provide a consistent improvement: while it can increase the MT-Bench score (+0.47 on Llama-2-7B-80k, +0.30 on Mixtral-8x22B-v0.1-4bit), it can also, in some cases, give worse results than the original Urial. Similarly, it yields mixed results when measured by win rate on AlpacaEval 2.0, which was not optimized by the greedy search. This analysis suggests that the most effective ICL demonstrations might vary across base LLMs, potentially because of differences in their pre-training. This could further explain why Urial underperforms on multiple models in Table 1, especially on those which have become available only recently and were not used for selecting the URIAL examples.
Table 4: Transferability of greedy search IC prompt to other LLMs. We use the 1st-turn score on MT-Bench and the length-controlled (LC) win rates on AlpacaEval (AE) 2.0 when using various IC prompts on different base models.
Model MT-B (1st) AE 2.0
Base model: Llama-2-7B-80k
Urial (3 examples) $5.19$ $1.81$
Urial + greedy search (1 ex.) $5.56$ $1.50$
Urial + greedy search (2 ex.) 5.57 $1.84$
Urial + greedy search (3 ex.) $5.10$ 2.91
Base model: Mixtral-8x22B-v0.1-4bit
Urial (3 examples) $8.28$ $14.77$
Urial + greedy search (1 ex.) $8.36$ 15.74
Urial + greedy search (2 ex.) $7.79$ $13.20$
Urial + greedy search (3 ex.) 8.58 $14.68$
5 Conclusions and discussion
In this work, we have first illustrated that ICL via Urial provides a good baseline for instruction-following alignment, but presents a few limitations: it mostly performs slightly worse than fine-tuning on single-turn conversations and does not generalize well to multi-round ones. Focusing on single-turn evaluation, we have shown that simply adding more in-context demonstrations is not sufficient to reliably improve performance, unlike what happens for less open-ended tasks (Agarwal et al., 2024; Bertsch et al., 2024). Even with more sophisticated approaches, fully closing the gap between ICL and IFT does not appear easy. We have provided several analyses of why this could be the case, showing that ICL for instruction following presents different behavior compared to the tasks discussed in Min et al. (2022): in our case, the demonstrations need to be carefully chosen and of high quality, with correct answers to each question. We conjecture that via ICL, the LLM can learn to infer the response style, but the overall learning signal is not sufficiently strong to benefit from a large amount of examples, despite the very long context windows which have become available only recently. Interestingly, a few incorrect ICL demonstrations are sufficient to severely degrade the performance.
In general, LLMs still require many different post-training steps beyond simple instruction following: domain-specific continued pre-training, long-context fine-tuning, fine-tuning for function calling, safety training against jailbreaks, prevention of prompt injections, etc. The difficulty of achieving even the baseline instruction-following alignment through ICL raises questions about the feasibility of obtaining more refined capabilities without fine-tuning.
Acknowledgements
We are grateful to OpenAI for providing us API credits and access to GPT-4-Base as part of the OpenAI Researcher Access Program. M.A. was supported by the Google Fellowship and Open Phil AI Fellowship.
Appendix A Experimental Details
In this section, we show more details about the experiments we conducted in the paper.
A.1 Datasets and base models
We select high-quality demonstrations for the base models to learn at test time from open-sourced instruction fine-tuning datasets: (a) Evol-Instruct-70k (Xu et al., 2024) contains 70k training examples with varying complexity and is well-known for the use to build the series of WizardLM models, and (b) OpenHermes-2.5 (Teknium, 2023) is a continuation of the OpenHermes-1 dataset, at a much larger scale, much more diverse, and much higher quality compilation. We employ Evol-Instruct-70k dataset, which leads to a better open-sourced aligned model in terms of the MT-Bench score, for the Random scaling experiment and use OpenHermes-2.5 for the Diversity scaling experiment since it contains annotations for the data source.
We compare ICL performance across multiple models that have sufficiently large context windows:
  1. 1.
    Mistral-7B-v0.2 (Jiang et al., 2023) is the newest 7B model from Mistral when we are writing and its instruct version significantly outperforms Mistral-7B-Instruct-v0.1 on MT-Bench. The trained context length of Mistral-7B-v0.2 is 32k tokens.
  2. 2.
    Llama-2-7B-80k (Fu et al., 2024) is a variant of Llama-2-7B model fine-tuned with 80k context on a carefully designed long-document data mixture.
  3. 3.
    Mixtral-8x22B-v0.1-4bit is a variant of Mixtral-8x22B-v0.1 (Jiang et al., 2024), which is a pre-trained generative sparse Mixture of Experts, quantized with 4-bit precision. It contains $\sim$176B parameters and $\sim$44B active during inference, and it has a 65k context window.
We use the same inference configuration as Urial (Lin et al., 2024). Concretely, we employ greedy decoding for all models, including base and instruction fine-tuned models, to maximize reproducibility and secure a fair and robust evaluation. Moreover, we use a repetition penalty on base models to prevent degeneration, 1.0 for MoE models and 1.15 for the remaining models.
A.2 Evaluation
MT-Bench (Zheng et al., 2023), consists of 80 high-quality and challenging questions that have multi-round interaction, designed to examine the multi-turn conversation and instruction-following capability of models. It features 8 common categories of user prompts: coding, math, reasoning, extraction, roleplay, writing, humanities/social science, and STEM.
AlpacaEval 2.0 (Li et al., 2023) provides 805 test instructions, on which we generate new responses using the target model, and then calculate the score by competing with the baseline model (i.e., GPT-4-Turbo) judged by a designated automatic evaluator. We apply the AlpacaEval 2.0 benchmark in our experiments to ensure that the effective demonstrations found through greedy search don’t overfit to MT-Bench.
A.3 Greedy search
We start with constructing a dataset that solely consists of high-quality examples from Evol-Instruct-70k. We then sample 100 high-quality examples as the initial search space. First, we employ GPT-3.5-Turbo as the judge to evaluate the quality of each data point in Evol-Instruct-70k with a prompt template as shown in Fig. 3. A thorough examination of the Evol-Instruct-70k dataset reveals a score distribution (Fig. 4), in which over 14k examples have a score of at least 4.5. Subsequently, we truncate the entire dataset by setting up a threshold of 4.5, resulting in a smaller dataset that is exclusively comprised of high-quality examples as indicated by GPT-3.5-Turbo.
Figure 3: The prompt template for evaluating the quality of data points using GPT-3.5-Turbo as the judge.
Next, we randomly sample 100 examples and then create 100 new prompts by adding each one of the sampled examples, as the fourth demonstration, to the original 3-shot Urial prompt. We evaluate the resulting instruction-following performance for each prompt using MT-Bench but with GPT-4-Turbo as the judge, and select the best fourth example based on the 1st-turn score. We continue our greedy search for the proper fifth and sixth demonstration in a restricted search space. Due to computational cost, we keep reducing the search space via applying a threshold, which is decided based on the 1st-turn MT-Bench score distribution we get in the previous round of searching.
A.4 Scaling experiments
The recently released LLMs with increasingly large context windows allow using many more shots for ICL than are typically used. Prior works (Agarwal et al., 2024; Bertsch et al., 2024) have shown that many-shot ICL, compared to few-shot ICL, can make task-specific fine-tuning less essential and allows LLMs to tackle a wider range of tasks without specification. Therefore, we want to examine if the instruction-following performance of base LLMs could benefit from many-shot ICL through various scaling experiments. In the following, we add more details about these scaling experiments, including the prompt template and data selection for ICL.
Figure 4: The distribution of data quality for Evol-Instruct-70k dataset as measured by using GPT-3.5-Turbo as the judge.
Figure 5: The prompt template for doing ICL in our work. Note that the words in orange are solely for illustration purposes and do not appear in the real prompts. The mixture of new demonstrations we test in our experiments is inserted before Urial demonstrations.
Random.
We construct the set of in-context demonstrations based on the high-quality dataset we create from Evol-Instruct-70k. Through randomly sampling from the high-quality dataset multiple times, we generate a series of in-context demonstration sets that each contains $N$ examples, where $N\in\{2,7,17,27,37,47\}$, and insert the $N$ demonstrations into the prompt template for ICL as shown in Fig 5. Note that the prompt template is strictly following the one used in Urial paper and we only replace the demonstrations used for the ICL purpose.
Greedy.
One could perceive the greedy search results of the proper fourth example as an alternative method of evaluating the data quality. An example adding to the context as the fourth demonstration results in a higher 1st-turn MT-Bench score, indicating its high quality in our ICL scenario. In this way, we can sort the 100 high-quality examples in descending order based on the score as evaluated by GPT-4-Turbo. Thus it allows us to create another series of in-context demonstration sets by selecting top-$N$ examples for each $N\in\{2,7,17,27,37,47\}$. The resulting mixture of in-context demonstrations is then placed in the corresponding location of the prompt template (Fig. 5).
Diversity.
Agarwal et al. (2024) suggest including diverse ICL examples could potentially lead to better generation, a stronger capability to handle more complex instructions and greater control over model outputs. Our work revisits this finding in the instruction-following task by scaling up examples from different sources. Concretely, we employ stratified sampling over the source annotation of OpenHermes-2.5 to collect $N$ diverse ICL examples, where $N\in\{3,5,10,20,30,40,50\}$. We get rid of the Urial demonstrations in this experiment since their source is unknown.
Figure 6: The prompt template for doing Unsupervised ICL in our work. Note that the words in orange are solely for illustration purposes and do not appear in the real prompts. The mixture of new demonstrations we test in our experiments is inserted before Urial demonstrations.
Unsupervised.
Motivated by encouraged results with unsupervised ICL in Agarwal et al. (2024), we examine the efficacy of scaling up the number of unsupervised queries. For simplicity, we reuse instructions from the experiment of scaling up diverse examples and remove corresponding answers. Specifically, the prompt template for unsupervised ICL scaling experiment consists of: (a) general rules that the model should follow when generating answers, (b) multiple instructions without answers, and (c) a few-shot prompt with ground truth answers for the desired output format and the Urial prompt is used in our work. We show the exact prompt template in Fig. 6.
Appendix B Additional experiments
The test set of MT-Bench (Zheng et al., 2023) is comprised of high-quality questions that belong to 8 common categories: coding, math, reasoning, extraction, roleplay, writing, humanities/social science, and STEM. In Table 5, we present the per-category performance on MT-Bench for each model in order to have a more detailed comparison between the Urial-aligned models and the corresponding instruct models. The results suggest fine-tuning is almost always better than Urial (ICL) and only underperforms in a minority of cases on particular base models, such as Mistral-7B-v0.1 and Llama-2-70B.
Table 5: A breakdown examination of Urial-aligned models and corresponding instruct models across different categories on MT-Bench. * denotes the results taken from the Urial GitHub repository.
Model coding math reasoning extraction humanities roleplay stem writing Average
Llama-2-7B + Urial * $1.65$ $1.60$ $3.45$ $3.40$ $8.08$ $7.48$ $6.80$ $6.20$ $4.83$
Llama-2-7B-Instruct 2.95 2.40 5.20 6.33 9.58 7.83 8.88 9.05 6.53
Llama-2-70B + Urial * 4.15 $3.60$ 6.10 7.70 $9.75$ $7.33$ $8.75$ 9.50 $7.11$
Llama-2-70B-Instruct $3.75$ 4.10 $5.95$ $7.40$ 9.85 7.90 9.13 9.50 7.20
Llama-3-8B + Urial * $4.15$ $2.60$ $3.50$ $5.25$ $8.90$ $7.30$ $8.15$ $6.13$ $5.75$
Llama-3-8B-Instruct 5.95 5.05 6.15 9.16 9.90 9.05 8.95 8.70 7.86
Llama-3-70B + Urial * $4.35$ $3.80$ $4.95$ $6.20$ $8.00$ $7.25$ $8.55$ $8.10$ $6.40$
Llama-3-70B-Instruct 7.85 7.35 6.25 9.75 10.00 9.30 9.60 9.80 8.74
Mistral-7B-v0.1 + Urial * 4.60 $3.40$ $4.90$ 7.75 $9.08$ 7.65 8.28 $7.75$ $6.67$
Mistral-7B-Instruct-v0.1 $4.35$ 3.95 6.30 $6.75$ 9.45 $7.45$ $7.70$ 8.85 6.85
Mistral-7B-v0.2 + Urial * $3.80$ $3.35$ $4.50$ $7.45$ $8.95$ $6.70$ $7.43$ $7.98$ $6.27$
Mistral-7B-Instruct-v0.2 5.45 3.40 6.50 8.50 9.90 8.65 9.30 9.40 7.64
Mixtral-8x22B-v0.1-4bit + Urial $5.25$ $5.85$ $6.00$ 9.20 9.80 $8.65$ $8.30$ $8.60$ $7.71$
Mixtral-8x22B-Instruct-v0.1-4bit 7.10 7.03 7.00 $9.10$ $9.65$ 8.90 9.65 9.70 8.52
GPT-4-Base + Urial $5.55$ $5.98$ $6.90$ $8.20$ $5.93$ $6.90$ $6.45$ $6.10$ $6.50$
GPT-4 (March 2023) * 8.55 6.80 9.00 9.38 9.95 8.90 9.70 9.65 8.99